OpenStudy (anonymous):

okay, I am having trouble with this will sumone explain it to me? Enter the quadratic equation in standard form. x^2 -5x-5=2x+3

Parth (parthkohli):

The standard form of a quadratic equation is basically: \(\Large \color{MidnightBlue}{\Rightarrow ax^2 + bx + c = 0 }\) Bring 0 on one side.

Parth (parthkohli):

From both sides, subtract 2x and 3 to bring it to the standard form.

OpenStudy (anonymous):

x^2 -5x-5=2x+3 x^2 -5x -5 -2x=2x+3-2x x^2 -7x-5=3 x^2 -7x -5 -3 = 3-3 x^2 -7x -8=0

OpenStudy (anmolsingh):

Well the standard form for a quadratic equation is Ax^2+Bx+C=0 where A,B,C are constants. For your problem, bring all the parts of the equation on one side and do the following, x^2 -5x-5=2x+3 x^2 -5x -5 -2x=2x+3-2x x^2 -7x-5=3 x^2 -7x -5 -3 = 3-3 x^2 -7x -8=0